A well-known brokerage firm executive claimed that 60% of investors are currently confident of meeting their investment goals. An XYZ Investor Optimism Survey, conducted over a two week period, found that in a sample of 200 people, 65% of them said they are confident of meeting their goals.

Test the claim that the proportion of people who are confident is larger than 60% at the 0.10 significance level.

The null and alternative hypothesis would be:

H0:μ=0.6H0:μ=0.6
H1:μ≠0.6H1:μ≠0.6

H0:μ=0.6H0:μ=0.6
H1:μ<0.6H1:μ<0.6

H0:μ=0.6H0:μ=0.6
H1:μ>0.6H1:μ>0.6

H0:p=0.6H0:p=0.6
H1:p≠0.6H1:p≠0.6

H0:p=0.6H0:p=0.6
H1:p>0.6H1:p>0.6

H0:p=0.6H0:p=0.6
H1:p<0.6H1:p<0.6



The test is:

two-tailed

left-tailed

right-tailed



The test statistic is:  (to 3 decimals)

The p-value is:  (to 4 decimals)

Based on this we:

  • Fail to reject the null hypothesis
  • Reject the null hypothesis
Answer #1

H0: p = 0.60

Ha: p > 0.60

The test is right tailed

Test statistics

z = ( - p) / sqrt [ p( 1 - p) / n ]

= (0.65 - 0.60) / sqrt ( 0.60 * ( 1 - 0.60) / 200 ]

= 1.443

p-value = P(Z > z)

= P(Z > 1.443)

= 1 - P(Z < 1.443)

= 1 - 0.9255

= 0.0745

Since p-value < 0.10 significance level, Reject the null hypothesis.